# Continuous Diffusion for Mixed-Type Tabular Data

### v3

@@ 4.2 Results
Table 1: Average performance rank of each generative model across eleven datasets. Per metric, bold indicates the best, underline the second best result. We assigned the rank 10 for CoDi on lending and diabetes, TabDDPM on acsincome and diabetes, SMOTE on acsincome and covertype.
@@
RMSE $\underline{3.400}_{\pm 3.382}$ $3.800_{\pm 2.482}$ $7.800_{\pm 1.470}$ $7.800_{\pm 2.227}$ $8.200_{\pm 1.470}$ $6.800_{\pm 1.939}$ $6.800_{\pm 1.720}$ $\mathbf{3.000}_{\pm 0.632}$ $\underline{3.400}_{\pm 2.059}$ $4.200_{\pm 2.227}$
F1 $3.667_{\pm 3.145}$ $6.333_{\pm 2.285}$ $8.333_{\pm 1.491}$ $8.000_{\pm 1.000}$ $4.167_{\pm 2.794}$ $6.667_{\pm 3.091}$ $7.500_{\pm 1.607}$ $3.833_{\pm 1.462}$ $\mathbf{3.167}_{\pm 1.344}$ $\underline{3.500}_{\pm 1.979}$
AUC $4.667_{\pm 2.749}$ $5.667_{\pm 2.055}$ $8.667_{\pm 1.106}$ $7.833_{\pm 1.067}$ $4.833_{\pm 2.794}$ $7.500_{\pm 2.872}$ $7.333_{\pm 1.374}$ $\underline{2.500}_{\pm 1.500}$ $\mathbf{2.333}_{\pm 0.943}$ $3.833_{\pm 1.675}$
L2 dist. of corr. $4.818_{\pm 2.918}$ $5.636_{\pm 1.872}$ $8.091_{\pm 1.781}$ $7.909_{\pm 1.564}$ $7.000_{\pm 3.191}$ $6.909_{\pm 2.429}$ $6.818_{\pm 1.402}$ $\underline{2.727}_{\pm 1.286}$ $\mathbf{2.273}_{\pm 0.862}$ $3.000_{\pm 1.595}$
Detection score $3.909_{\pm 3.502}$ $6.182_{\pm 1.696}$ $8.818_{\pm 1.466}$ $7.273_{\pm 1.213}$ $5.000_{\pm 3.045}$ $8.091_{\pm 2.391}$ $6.000_{\pm 1.595}$ $3.909_{\pm 1.164}$ $\mathbf{2.455}_{\pm 1.827}$ $\underline{3.545}_{\pm 1.725}$
JSD $7.182_{\pm 2.167}$ $\mathbf{1.273}_{\pm 0.617}$ $8.182_{\pm 1.641}$ $8.818_{\pm 1.029}$ $6.909_{\pm 2.314}$ $7.000_{\pm 1.651}$ $6.545_{\pm 1.305}$ $\underline{2.455}_{\pm 1.076}$ $3.091_{\pm 0.793}$ $3.727_{\pm 1.052}$
WD $\underline{3.091}_{\pm 3.315}$ $5.636_{\pm 1.611}$ $7.545_{\pm 1.827}$ $8.000_{\pm 1.477}$ $6.455_{\pm 3.144}$ $8.364_{\pm 1.823}$ $5.727_{\pm 2.339}$ $4.182_{\pm 1.466}$ $3.182_{\pm 1.192}$ $\mathbf{3.000}_{\pm 1.954}$
DCR $6.000_{\pm 2.558}$ $6.182_{\pm 2.328}$ $8.455_{\pm 1.725}$ $6.182_{\pm 3.186}$ $4.455_{\pm 3.726}$ $6.545_{\pm 2.426}$ $5.909_{\pm 1.676}$ $4.091_{\pm 2.678}$ $\underline{3.818}_{\pm 2.124}$ $\mathbf{3.545}_{\pm 1.924}$
Table 1 shows the average rank of each generative model across all datasets for the considered metrics. The ranks in terms of the F1 and AUC scores are averaged over the classification task datasets. Likewise, the RMSE rank averages include the regression task datasets. We assign the maximum possible rank when a model could not be trained on a given dataset or could not be evaluated in reasonable time. This includes TabDDPM, which outputs NaNs for acsincome and diabetes and CoDi, which we consider to be prohibitively expensive to train on diabetes (estimated 14.5 hours) and lending (estimated 60 hours). Similarly, SMOTE is very inefficient in sampling for large datasets (78 min for 1000 samples on acsincome and 182 min on covertype) and does not finish the evaluation within 12 hours. The dataset-specific results (including standard errors) and average metrics over all datasets are detailed in Appendix R. We provide visualizations of the captured correlations in the synthetic sample compared to the real training set in Appendix Q and distribution plots for a qualitative comparison in Appendix P.
@@
CDTD consistently outperforms the considered benchmark models in most sample quality metrics. Specifically, we see a major performance edge in terms of the detection score, the L2 distance of the correlation matrices and the regression-based metrics. Only for the Jensen-Shannon divergence ARF, a tree-based method that is expected to model categorical features particularly well, outperforms CDTD. Interestingly, CDTD performs similar to TabDDPM on F1 scores, but outperforms it dramatically for regression tasks. TabDDPM appears to favor modeling categorical features accurately, thereby sacrificing continuous features, as visualized in Appendix Q. TabSyn, a latent-space diffusion model, performs worse than CDTD and often TabDDPM, which define diffusion in data space. In Appendix M, we further compare CDTD and TabSyn and investigate the benefits of defining a diffusion model in data space. By utilizing score interpolation, CDTD is able to model intricate correlation structure more accurately than other frameworks. Most importantly, type-specific noise schedules mostly outperform the feature-specific and single noise schedule variants. This illustrates the importance of accounting for the high heterogeneity in tabular data on the feature type level. The different noise schedules per feature, however, appear to force too many constraints on the model and thus, decrease sample quality. Per-feature noise schedules would require more training steps to converge, as can be seen in Appendix O.
Table 2: Ablation study for five CDTD configurations with progressive addition of model components. We report the median performance metrics over acsincome, adult, beijing and churn.
Config. A B C D CDTD (per type)
RMSE (abs. diff.; $\downarrow$) 0.041 0.042 0.043 0.037 0.033
F1 (abs. diff.; $\downarrow$) 0.012 0.013 0.012 0.016 0.015
AUC (abs. diff.; $\downarrow$) 0.004 0.005 0.005 0.004 0.004
L2 distance of corr.  ($\downarrow$) 0.131 0.124 0.146 0.118 0.127
Detection score ($\downarrow$) 0.577 0.583 0.590 0.561 0.560
JSD ($\downarrow$) 0.011 0.011 0.011 0.012 0.013
WD ($\downarrow$) 0.004 0.005 0.005 0.003 0.003
DCR (abs. diff. to test; $\downarrow$) 0.405 0.361 0.386 0.299 0.372
Figure 3: Average training and sampling wall-clock time for 1000 samples (excl. acsincome, diabetes, lending).
@@
Figure 3 shows the average wall-clock time over all (for all models feasible) datasets for training as well as the time for sampling 1000 data points for each baseline model and the per feature CDTD variant (see Appendix T for details). We exclude SMOTE due to its considerably longer sampling with an average of 1377 seconds for 1000 samples. CDTD’s use of embeddings (instead of one-hot encoding) for categorical features drastically reduces training times and thus, improves scaling to increasing number of categories. The ODE formulation of the diffusion process implies competitive sampling speeds, in particular compared to the diffusion-based benchmarks CoDi, TabDDPM and TabSyn. Despite TabSyn utilizing a separately trained encoder, this does not result in a lower-dimensional latent space and therefore, does not speed up sampling.
@@
We conduct an ablation study over four datasets to investigate the separate components of our CDTD framework. The results are given in Table 2 (detailed results are in Appendix S). The baseline model Config. A includes a single noise schedule with the original piece-wise linear formulation (Dieleman et al., 2022) without loss normalization, improved model initialization or adaptive normalization, and the CE and MSE losses are naively averaged. Note that this configuration still is a novel contribution to the literature. Config. B adds our feature homogenization (i.e., loss normalization, improved initialization and adaptive normalization schemes), Config. C adds our proposed functional form for a single noise schedule with uniform initialization, and Config. D imposes per-type noise schedules. Lastly, we add the suggested (low noise level) overweighting timewarping initialization to arrive at the full CDTD (per type) model. We see the switch from the piece-wise linear functional form to our more robust noise schedule variant slightly harms sample quality. However, the per-type variant and the more informed initialization scheme compensate for this difference. Main differences are in the RMSE and detection score as well as training efficiency (the loss calibration and improved initialization facilitate model convergence). The final model especially works well on the larger datasets compared to the baseline (see Appendix S), as smaller datasets are relatively easy to fit with 3 million parameters, even without any model improvements. We investigate the sensitivity of CDTD to important hyperparameters in Appendix L.
@@ 5 Conclusion and Discussion
We propose a Continuous Diffusion model for mixed-type Tabular Data (CDTD) that combines score matching and score interpolation and imposes Gaussian diffusion processes on both continuous and embedded categorical features. We compared CDTD to various benchmark models and to a single noise schedule as typically used in image diffusion models. Our results indicate that addressing the high feature heterogeneity in tabular data on the feature type level and aligning type-specific diffusion elements, such as the noise schedules or losses, substantially benefits sample quality. Moreover, CDTD shows vastly improved scalability and can accommodate an arbitrary number of categories.
Our paper serves as an important step to customizing the diffusion probabilistic framework to tabular data. In particular, the common type of noise schedules allows for an easy to extend framework that might accelerate progress on diffusion models for tabular data. Crucially, CDTD allows the direct application of diffusion-related advances from the image domain, like classifier-free guidance, to tabular data without the need for a latent encoding. We leave further extensions to the tabular data domain, e.g., the exploration of accelerated sampling, efficient score model architectures, different forms of adaptive noise schedules, or the adaption to the data imputation task for future work.
Finally, we want to emphasize the potential misuse of synthetic data to support unwarranted claims. Any generated data should therefore not be blindly trusted, and synthetic data based inferences should always be compared to results from the real data. However, the correct use of generative models enables better privacy preservation and facilitates data sharing and open science practices.
@@
The main limitation of CDTD is the addition of hyperparameters, and tuning hyperparameters of a generative model can be a costly endeavor. However, our results also show that (1) a per type schedule is most often optimal and (2) our default hyperparameters perform well across a diverse set of datasets. Dieleman et al. (2022) show that the results of score interpolation for text data can be sensitive to the initialization of the embeddings. We have not encountered similar problems on tabular datasets (see Table 7). While the DCR indicates no privacy issues for the benchmark datasets used, additional caution must be taken when generating synthetic data from privacy sensitive sources. Lastly, for specific types of tabular data, such as time-series, our model may be outperformed by other generative models specialized for that type. While CDTD could be directly used for imputation using RePaint (Lugmayr et al., 2022), a separate training process is required to achieve the best results (Liu et al., 2024). Therefore, we leave the adaption of CDTD to the imputation task for future work.
@@ Acknowledgements
This work used the Dutch national e-infrastructure with the support of the SURF Cooperative using grant no. EINF-7437. We would also like to thank Sander Dieleman for helpful discussions.
Appendix A Loss Calibration
A priori, we let the model be indifferent between features, that is, we scale the loss of each feature such that at the terminal timestep the same loss is attained. Here, the signal-to-noise ratio is sufficiently low to approximate a situation in which the model has no information about the data. Thus, we are looking for calibrated losses $\mathcal{L}_{\text{MSE}}^{\ast}(x_{\text{cont}}^{(i)},1)$ and $\mathcal{L}_{\text{CE}}^{\ast}(x_{\text{cat}}^{(j)},1)$ which at $t=1$ achieve unit loss in expectation.
@@ Appendix A Loss Calibration
where $F_{\bm{\theta}}^{(i)}$ denotes the neural network output for feature $i$ that parameterizes the score model $s_{\bm{\theta}}$. The parameters $c_{\text{skip}}(t)=\sigma_{\text{data}}^{2}/(\sigma^{2}(t)+\sigma_{\text{data}}^{2})$ and $c_{\text{out}}(t)=\sigma(t)\cdot\sigma_{\text{data}}/(\sqrt{\sigma^{2}(t)+\sigma_{\text{data}}^{2}})$ depend on $\sigma(t)$ (and $\sigma_{\text{data}}$) and therefore on timestep $t$. For $t\rightarrow 1$, $\sigma(t)$ approaches the maximum noise level $\sigma_{\text{cont},\max}$ and $c_{\text{skip}}(t)\rightarrow 0$ and $c_{\text{out}}(t)\rightarrow 1$ such that the score model directly predicts the data at high noise levels. For $t\rightarrow 0$, the model shifts increasingly towards predicting the error that has been added to the true data. In the EDM parameterization, the explicit timestep weight (used to achieve a unit loss across timesteps at initialization, see Appendix B) is $\lambda(t)=1/c_{\text{out}}(t)^{2}\approx 1$ for $t=1$.
At the terminal timestep $t=1$, we now have:
@@
For a single categorical feature, $x^{(j)}_{\text{cat}}$ is distributed according to the proportions $p_{c}$ (for categories $c=1,\dots,C$). The denoising model for score interpolation is trained with the CE loss:
$\mathcal{L}_{\text{CE}}(x^{(j)}_{\text{cat}},t)=-\sum_{c=1}^{C}I(x^{(j)}_{\text{cat}}=c)\log F_{{\bm{\theta}},c}^{(j)}\ ,$
where $F_{{\bm{\theta}},c}^{(j)}$ denotes the score model’s prediction of the class probability at timestep $t$. Without information, it is optimal to assign the $c$-th category the same proportion as in the training set. At $t=1$, we thus let $F_{{\bm{\theta}},c}^{(j)}=p_{c}$ such that the minimum loss equals:
$\mathbb{E}_{p(x^{(j)}_{\text{cat}})}[\mathcal{L}_{\text{CE}}(x^{(j)}_{\text{cat}},1)]$ $=-\mathbb{E}_{p(x^{(j)}_{\text{cat}})}\sum_{c=1}^{C}I(x^{(j)}_{\text{cat}}=c)\log F_{{\bm{\theta}},c}^{(j)}\ ,$ (9)
$=-\sum_{c=1}^{C}\mathbb{E}_{p(x^{(j)}_{\text{cat}})}[I(x^{(j)}_{\text{cat}}=c)\log p_{c}]\ ,$ (10)
$=-\sum_{c=1}^{C}p_{c}\log p_{c}.$ (11)
We use the training set proportions to compute the normalization constant $Z_{j}=\leavevmode\nobreak\ -\sum_{c=1}^{C}p_{c}\log p_{c}$ to calibrate the loss for categorical features. Then,
@@
Appendix B Output Layer Initialization
At initialization, we want the neural network to reflect the state of no information (see Appendix A). Likewise, our goal is a loss of one across all features and timesteps.
For continuous features $i$, we initialize the output layer weights (and biases) to zero such that the output of the score model for a single continuous feature, $F_{\bm{\theta}}^{(i)}$, is also zero. Since we use the EDM parameterization (Karras et al., 2022), we apply the associated explicit timestep weight $\lambda(t)=\frac{\sigma^{2}(t)+\sigma_{\text{data}}^{2}}{(\sigma(t)\cdot\sigma_{\text{data}})^{2}}$. This is explicitly designed to achieve a unit loss across timesteps at initialization and we show this analytically below. We denote the variances of the data $x_{\text{cont}}^{(i)}$ and of the Gaussian noise $\epsilon$ at time $t$ as $\sigma_{\text{data}}^{2}$ and $\sigma^{2}(t)$, respectively.

### v4

@@ 4.2 Results
Table 1: Average performance rank of each generative model across eleven datasets. Per metric, bold indicates the best, underline the second best result. We assigned the rank 10 for CoDi on lending and diabetes, TabDDPM on acsincome and diabetes, SMOTE on acsincome and covertype. RMSE, F1, AUC and DCR are measured in abs. differences to the real test set.
@@
RMSE $3.6_{\pm 3.3}$ $3.4_{\pm 2.1}$ $8.0_{\pm 1.6}$ $7.8_{\pm 2.2}$ $8.4_{\pm 1.0}$ $7.0_{\pm 1.8}$ $7.2_{\pm 1.5}$ $4.0_{\pm 1.7}$ $\mathbf{2.6}_{\pm 1.0}$ $\underline{3.2}_{\pm 1.5}$
F1 $4.2_{\pm 2.9}$ $6.2_{\pm 2.3}$ $8.0_{\pm 1.4}$ $8.0_{\pm 1.0}$ $4.3_{\pm 3.1}$ $6.5_{\pm 3.0}$ $8.2_{\pm 1.7}$ $4.0_{\pm 1.1}$ $\mathbf{2.3}_{\pm 1.4}$ $\underline{3.5}_{\pm 1.4}$
AUC $4.3_{\pm 3.0}$ $5.7_{\pm 1.9}$ $8.3_{\pm 1.2}$ $7.8_{\pm 1.1}$ $4.3_{\pm 3.1}$ $6.8_{\pm 3.1}$ $8.2_{\pm 1.6}$ $\underline{3.3}_{\pm 1.1}$ $\mathbf{2.7}_{\pm 1.5}$ $3.7_{\pm 1.2}$
L2 dist. of corr. $5.0_{\pm 2.8}$ $5.7_{\pm 2.2}$ $8.3_{\pm 1.9}$ $7.9_{\pm 1.4}$ $6.0_{\pm 3.3}$ $7.0_{\pm 2.5}$ $7.1_{\pm 1.2}$ $3.4_{\pm 1.1}$ $\mathbf{2.0}_{\pm 0.8}$ $\underline{2.8}_{\pm 1.5}$
Detection score $5.7_{\pm 2.5}$ $6.1_{\pm 1.7}$ $8.8_{\pm 1.5}$ $7.6_{\pm 1.1}$ $4.8_{\pm 3.2}$ $7.9_{\pm 2.4}$ $6.2_{\pm 1.6}$ $\underline{3.1}_{\pm 1.5}$ $\mathbf{1.6}_{\pm 1.0}$ $3.3_{\pm 1.1}$
JSD $7.2_{\pm 2.2}$ $\mathbf{1.2}_{\pm 0.4}$ $7.6_{\pm 2.4}$ $8.8_{\pm 1.0}$ $6.9_{\pm 2.1}$ $7.2_{\pm 1.5}$ $6.8_{\pm 1.3}$ $\underline{2.5}_{\pm 0.7}$ $2.8_{\pm 1.1}$ $4.1_{\pm 0.9}$
WD $\underline{3.2}_{\pm 3.3}$ $5.8_{\pm 1.8}$ $7.4_{\pm 2.4}$ $7.9_{\pm 1.6}$ $5.6_{\pm 3.5}$ $8.4_{\pm 1.8}$ $5.8_{\pm 2.0}$ $4.5_{\pm 1.6}$ $\mathbf{3.2}_{\pm 1.5}$ $3.4_{\pm 1.9}$
DCR $5.8_{\pm 2.7}$ $6.5_{\pm 2.1}$ $8.4_{\pm 1.7}$ $6.1_{\pm 2.9}$ $4.5_{\pm 3.4}$ $6.6_{\pm 2.4}$ $6.3_{\pm 2.0}$ $4.1_{\pm 2.5}$ $\underline{3.9}_{\pm 2.3}$ $\mathbf{3.0}_{\pm 2.0}$
Table 1 shows the average rank of each generative model across all datasets for the considered metrics. Detailed results (including standard errors) for each model and dataset are reported in Appendix S. The ranks in terms of the F1 and AUC scores are averaged over the classification task datasets. Likewise, the RMSE rank averages include the regression task datasets. We assign the maximum possible rank when a model could not be trained on a given dataset or could not be evaluated in reasonable time. This includes TabDDPM, which outputs NaNs for acsincome and diabetes and CoDi, which we consider to be prohibitively expensive to train on diabetes (estimated 14.5 hours) and lending (estimated 60 hours). Similarly, SMOTE is very inefficient in sampling for large datasets (78 min for 1000 samples on acsincome and 182 min on covertype) and does not finish the evaluation within 12 hours. We provide visualizations of the captured correlations in the synthetic sample compared to the real training set in Appendix R and distribution plots for a qualitative comparison in Appendix Q.
@@
CDTD consistently outperforms the considered benchmark models in most sample quality metrics. Specifically, we see a major performance edge in terms of the detection score, the L2 distance of the correlation matrices and the ML efficiency metrics. Using score interpolation, CDTD is able to model the intricate correlation structure more accurately than other frameworks. Due to their inductive biases, SMOTE and ARF perform very well in terms of univariate distribution fit for continuous features (as measured by WD) and categorical (as measured by JSD), respectively. However, in both of those cases, CDTD performs competitively, in particular when compared to the diffusion-based models. Interestingly, TabSyn, a latent space diffusion model, performs considerably worse than CDTD and often TabDDPM, which define diffusion in data space. In Appendix N, we further compare CDTD and TabSyn and investigate the benefits of defining a diffusion model in data space.
Most importantly, type-specific noise schedules mostly outperform the feature-specific and single noise schedule variants. This illustrates the necessity to account for the high heterogeneity in tabular data on the feature-type level. Having distinct noise schedules per feature instead appears to force too many constraints on the model and, thus, decreases sample quality. Per-feature noise schedules would also require more training steps to fully converge, as can be seen in Appendix P. We investigate the sensitivity of CDTD to important hyperparameters in Appendix M.
Table 2: Ablation study for five CDTD configurations. We report the median performance. The grey column depicts the impact of our data type homogenization.
{NiceTabular}
[…]
RMSE (abs. diff.; $\downarrow$) 0.090 0.069 0.060 0.058 0.055
F1 (abs. diff.; $\downarrow$) 0.015 0.011 0.010 0.006 0.009
AUC (abs. diff.; $\downarrow$) 0.008 0.005 0.005 0.005 0.006
L2 distance of corr.  ($\downarrow$) 0.483 0.421 0.457 0.450 0.444
Detection score ($\downarrow$) 0.762 0.739 0.774 0.662 0.701
JSD ($\downarrow$) 0.010 0.013 0.011 0.011 0.011
WD ($\downarrow$) 0.008 0.006 0.007 0.005 0.006
DCR (abs. diff. to test; $\downarrow$) 0.639 0.552 0.574 0.350 0.544
Figure 3: Average training and sampling wall-clock times (excl. SMOTE, acsincome, diabetes, lending).
@@
Figure 3 shows the average training and sampling wall-clock times (for all feasible models) over all datasets (see Appendix U for details on sample quality as a function of sampling time). We exclude SMOTE due to its considerably longer sampling times with an average of 1377 seconds for 1000 samples. CDTD’s use of embeddings (instead of one-hot encoding) for categorical features improves its scaling to increasing number of categories and thus, drastically reduces training times. The sampling speed of CDTD is competitive, in particular compared to the diffusion-based benchmarks CoDi, TabDDPM and TabSyn.
@@
We investigate the separate components of our CDTD framework. The summarized results are given in Table 4.2 and the detailed results in Appendix T. The baseline model Config. A includes a single noise schedule with the original piece-wise linear functional form (Dieleman et al., 2022) and the CE and MSE losses are naively averaged. Note that this configuration is still a novel contribution. Config. B adds our data type homogenization (i.e., loss calibration, improved initialization and time-dependent normalization schemes), Config. C adds our proposed functional form for a single noise schedule with uniform initialization, and Config. D imposes per-type noise schedules. Lastly, we increase the importance of low noise levels at initialization to arrive at the full CDTD (per type) model.
The results show that data type homogenization benefits sample quality significantly. Metrics associated with continuous features, i.e., RMSE and WD, as well as those relying on all features being generated well, i.e., the detection score, L2 distance of corr. matrices and DCR, improve dramatically. Switching from the piecewise linear noise schedule to our more robust functional form slightly harms sample quality. However, the per-type variant and the improved initialization more than compensate for this. The latter appears to trade-off some sample quality for faster convergence during training.
@@ 5 Conclusion and Discussion
In this paper, we introduce a Continuous Diffusion model for mixed-type Tabular Data (CDTD) that combines score matching and score interpolation and imposes Gaussian diffusion processes on both continuous and embedded categorical features. Our results indicate that addressing the high feature heterogeneity in tabular data on the feature type level by aligning type-specific diffusion elements, such as the noise schedules or losses, substantially benefits sample quality. Moreover, CDTD shows vastly improved scalability and can accommodate an arbitrary number of categories.
Our paper serves as an important step to customizing score-based models to tabular data. The common type of noise schedules allows for an easy to extend framework. Crucially, CDTD allows the direct application of diffusion-related advances from the image domain, like classifier-free guidance, to tabular data without the need for a latent encoding. We leave further extensions to the tabular data domain, e.g., the exploration of accelerated sampling, efficient score model architectures, or the adaptation to the data imputation task for future work.
Finally, we want to warn against the potential misuse of synthetic data to support unwarranted claims. Any generated data should not be blindly trusted, and synthetic data based inferences should always be compared to results from the real data. However, the correct use of generative models for tabular data enables better privacy preservation and facilitates data sharing and open science practices.
@@
The main limitation of CDTD is the addition of hyperparameters, and tuning hyperparameters of a generative model can be a costly endeavor. However, our results also show that (1) a per type schedule is most often optimal and (2) our default hyperparameters perform well across a diverse set of datasets. Dieleman et al. (2022) show that the results of score interpolation for text data can be sensitive to the initialization of the embeddings. We have not encountered similar problems on tabular datasets (see Table 7). While the DCR indicates no privacy issues for the benchmark datasets used, additional caution must be taken when generating synthetic data from privacy sensitive sources. Lastly, for specific types of tabular data, such as time-series, our model may be outperformed by other generative models specialized for that type. While CDTD could be directly used for imputation using RePaint (Lugmayr et al., 2022), a separate training process is required to achieve the best results (Liu et al., 2024). Therefore, we leave the adaptation of CDTD to the imputation task for future work.
@@ Acknowledgements
This work used the Dutch national e-infrastructure with the support of the SURF Cooperative using grant no. EINF-7437. We also thank Sander Dieleman for helpful discussions.
Appendix A Noise implications of heterogeneous cardinalities
One distinct characteristic of mixed-type tabular data is that each categorical feature may have a different cardinality, i.e., a different number of categories. Below, we briefly empirically investigate how this data characteristic warrants a rethinking of diffusion noise schedules. We show that, under some assumptions, categorical features of higher cardinality require more added noise than lower cardinality features to remove the same amount of signal.
Let $x_{\text{cat}}^{(j)}$ be a single categorical feature with $C_{j}$ categories. We assume that each category $c$ has equal probability $\mathrm{P}(c)=1/C_{j}$ and train a CDTD model with a fixed, linear noise schedule to learn the distribution of a single $x_{\text{cat}}^{(j)}$. For each class $c\in\{1,\dots,C_{j}\}$, we extract the associated learned embedding ${\mathbf{e}}_{c}^{(j)}$. For 500 timesteps $t$, linearly spaced on $[0,1]$, we derive the associated noise levels $\sigma(t)$ from the learned noise schedules and for each $\sigma(t)$, we sample 1000 noisy embeddings ${\mathbf{e}}_{c}^{(j)}(t)={\mathbf{e}}_{c}^{(j)}+\sigma(t){\mathbf{\epsilon}}$ such that ${\mathbf{\epsilon}}\sim\mathcal{N}({\bm{0}},{\mathbf{I}})$. We input all ${\mathbf{e}}_{c}^{(j)}(t)$ into the score model to compute the average (calibrated) cross-entropy (CE) loss. We define the remaining signal of the feature at a given $t$ as $S_{c}=1-\text{CE}_{c}(t)$, where $\text{CE}_{c}$ reflects the CE loss when using the noisy embedding for class $c$ to predict the true class. Lastly, we compute the overall remaining signal over all classes as $\max_{c}S_{c}$. Figure 4 shows the result of repeating this procedure with varying $C_{j}$ and four-dimensional embeddings. As we can see, lower cardinality features, i.e., those with a low $C_{j}$, attain the same signal only at a higher noise levels $\sigma(t)$. This example of the effect of feature heterogeneity in tabular data encourages us to investigate feature-specific noise schedules.
Figure 4: Comparison of signal remaining in embedded categorical features given by $\max_{c}S_{c}$, where $S_{c}=1-\text{CE}_{c}(t)$ and $\text{CE}_{c}$ reflects the CE loss when using the noisy embedding for class $c\in\{1,\dots,C_{j}\}$ to predict the true class. Lower cardinality features tend to require a higher noise level $\sigma(t)$ to achieve the same amount of remaining signal as higher cardinality features.
Appendix B Loss Calibration
Under Assumption 1, the signal-to-noise ratio at the terminal timestep is sufficiently low to approximate a situation in which the model has no information about the data. We want to let the model be indifferent between features, that is, we scale the loss of each feature such that at the terminal timestep the same expected loss is attained. Therefore, we are looking for scaled losses $\mathcal{L}_{\text{MSE}}^{\ast}(x_{\text{cont}}^{(i)},1)$ and $\mathcal{L}_{\text{CE}}^{\ast}(x_{\text{cat}}^{(j)},1)$ which at $t=1$ achieve unit loss in expectation.
@@ Appendix B Loss Calibration
where $F_{\bm{\theta}}^{(i)}$ denotes the neural network output for feature $i$ that parameterizes the score model $s_{\bm{\theta}}$. The parameters $c_{\text{skip}}(t)=\sigma_{\text{data}}^{2}/(\sigma^{2}(t)+\sigma_{\text{data}}^{2})$ and $c_{\text{out}}(t)=\sigma(t)\cdot\sigma_{\text{data}}/(\sqrt{\sigma^{2}(t)+\sigma_{\text{data}}^{2}})$ depend on $\sigma(t)$ (and $\sigma_{\text{data}}$) and therefore on timestep $t$. For $t\rightarrow 1$, $\sigma(t)$ approaches the maximum noise level $\sigma_{\text{cont},\max}$ and $c_{\text{skip}}(t)\rightarrow 0$ and $c_{\text{out}}(t)\rightarrow 1$ such that the score model directly predicts the data at high noise levels. For $t\rightarrow 0$, the model shifts increasingly towards predicting the error that has been added to the true data. In the EDM parameterization, the explicit timestep weight (used to achieve a unit loss across timesteps at initialization, see Appendix C) is $\lambda(t)=1/c_{\text{out}}(t)^{2}\approx 1$ for $t=1$.
At the terminal timestep $t=1$, we then have:
@@
For a single categorical feature, $x^{(j)}_{\text{cat}}$ is distributed according to the proportions $p_{c}$ (for categories $c=1,\dots,C_{j}$). The denoising model for score interpolation is trained with the CE loss:
$\mathcal{L}_{\text{CE}}(x^{(j)}_{\text{cat}},t)=-\sum_{c=1}^{C_{j}}I(x^{(j)}_{\text{cat}}=c)\log F_{{\bm{\theta}},c}^{(j)}\ ,$
where $F_{{\bm{\theta}},c}^{(j)}$ denotes the score model’s prediction of the class probability at timestep $t$. Without information, it is optimal to assign the $c$-th category the same proportion as in the training set. At $t=1$, we thus let $F_{{\bm{\theta}},c}^{(j)}=p_{c}$ such that the loss equals:
$\mathbb{E}_{p(x^{(j)}_{\text{cat}})}[\mathcal{L}_{\text{CE}}(x^{(j)}_{\text{cat}},1)]$ $=-\mathbb{E}_{p(x^{(j)}_{\text{cat}})}\sum_{c=1}^{C_{j}}I(x^{(j)}_{\text{cat}}=c)\log F_{{\bm{\theta}},c}^{(j)}\ ,$ (8)
$=-\sum_{c=1}^{C_{j}}\mathbb{E}_{p(x^{(j)}_{\text{cat}})}[I(x^{(j)}_{\text{cat}}=c)\log p_{c}]\ ,$ (9)
$=-\sum_{c=1}^{C_{j}}p_{c}\log p_{c}.$ (10)
Note that this is the feature-specific entropy and we can use the training set proportions to compute this normalization constant $Z_{j}=\leavevmode\nobreak\ -\sum_{c=1}^{C_{j}}p_{c}\log p_{c}$ to scale the loss for categorical features. Then,
@@
Appendix C Output Layer Initialization
At initialization, we want the neural network to reflect the state of no information (see Appendix B). Likewise, our goal is a loss of one across all features and timesteps.
For continuous features $i$, we initialize the output layer weights (and biases) to zero such that the output of the score model for a single continuous feature, $F_{\bm{\theta}}^{(i)}$, is also zero. Since we use the EDM parameterization (Karras et al., 2022), we apply the associated explicit timestep weight $\lambda(t)=\frac{\sigma^{2}(t)+\sigma_{\text{data}}^{2}}{(\sigma(t)\cdot\sigma_{\text{data}})^{2}}$. This is explicitly designed to achieve a unit loss across timesteps at initialization and we show this analytically below. We denote the variances of the data $x_{\text{cont}}^{(i)}$ and of the Gaussian noise $\epsilon$ at time $t$ as $\sigma_{\text{data}}^{2}$ and $\sigma^{2}(t)$, respectively. At initialization we have: